Directions: find the measure of the given angles.
An angle whose supplement is 3 times the measure of its complement.

Found 2 solutions by venugopalramana, Ms Ladii:
Answer by venugopalramana(3286)

let the angle =x
its complement =90-x
three times the complement =3*(90-x)=270-3x
its supplement = 180-x
so
180-x=270-3x
3x-x=270-180=90
2x=90
x=90/2=45
so the angle is 45 degrees

Answer by Ms Ladii(3)

1.) WRITE A LET STATMENT:
LET THE ANGLE =x
ITS COMPLEMENT =90-x
THREE TIMES ITS COMPLEMENT =3*(90-x)=270-3x
ITS SUPPLEMENT = 180-x
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
180-x=270-3x
3x-x=270-180=90
2x=90
x=90/2=45
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
THEREFORE THE ANGLE IS 45 DEGREES